OpenStudy (anonymous):

Find the circumference of a circle with radius 14 m

OpenStudy (anonymous):

circumference = 2*pi*r=2*3.14*14=2* 22/7*14

OpenStudy (anonymous):

\[2*\frac{ 22 }{ 7 }*14\]

OpenStudy (anonymous):

88.. right?

OpenStudy (anonymous):

correct ;)

OpenStudy (anonymous):

awesome!! thanks :)

OpenStudy (anonymous):

;) Yo'r welcome